# Explainable Hyperdimensional Computing for Balancing Privacy and Transparency in Additive Manufacturing Monitoring

### v1

@@ 6 Conclusion and Future Works
This paper introduced the Differential Privacy-HyperDimensional computing (DP-HD) framework as an innovative solution for addressing privacy concerns in Additive Manufacturing (AM). By integrating Differential Privacy (DP) mechanisms within the HyperDimensional computing (HD) model, the DP-HD framework ensures the protection of sensitive data during training and inference while maintaining high model performance. The key advantage of DP-HD is its explainability, enabling the prediction of noise impact on model accuracy, which is critical for in-situ sensing applications in AM. Our experiments demonstrated that DP-HD preserves privacy without significantly compromising accuracy, showcasing its robustness and suitability for real-time monitoring in AM. The ability to balance privacy and performance through explainable insights sets DP-HD apart, making it a valuable tool for enhancing manufacturing precision while safeguarding data privacy. While the DP-HD framework demonstrates significant promise, certain aspects warrant further exploration to enhance its applicability. Currently, finding the optimal encoding strategy relies on selecting training and validation subsets. Future work focuses on developing mathematical methods to automatically determine the best standard deviation for generating random basis vectors based on data distribution. Additionally, due to DP-HD’s success in privacy-preserving techniques, the framework can be integrated within a federated learning structure to leverage distributed data processing while maintaining robust privacy measures.
@@ Acknowledgment
This work was supported by the National Science Foundation [grant numbers 2127780, 2312517]; the Semiconductor Research Corporation (SRC); the Office of Naval Research [grant numbers N00014-21-1-2225, N00014-22-1-2067]; the Air Force Office of Scientific Research [grant number FA9550-22-1-0253]; UConn Startup Funding, and generous gifts from Xilinx and Cisco. The authors gratefully acknowledged the valuable contributions from the National Institute of Standards and Technology (NIST), in particular, Dr. Brandon Lane, for sharing data for this research.

### v3

@@ 5.2 Inference Privacy
Figure 11: Decoding optical images of laser powder bed fusion process to capture the confidential information.
@@
6 Discussion
In this paper, we introduced the DP-HD framework to secure sensitive information from in-situ sensing in AM environments. By combining DP with HD, DP-HD enhances privacy protection across both training and inference phases while achieving a balance between accuracy and privacy. This balance is quantified using an SNR, which measures the contribution of training samples to HD performance relative to the noise from DP. Below, we discuss the core components of our framework, as well as its strengths and limitations.
Our DP-HD framework leverages HD encoding to map feature vectors into a high-dimensional space, using a mathematically grounded approach that relies on principles inspired by Bochner’s theorem [80]. According to Bochner’s theorem, any continuous, shift-invariant kernel can be represented through a Fourier transform. Using Random Fourier Features (RFF) [81], our framework approximates this mapping, which allows the model to generalize across distributions effectively. Rather than depending on specific HD parameters or sample size, the optimal variance for RFF is derived directly from the underlying data distribution [82], allowing our model to adapt well to datasets with broad and representative distributions. Consequently, this encoding technique ensures that conclusions drawn from one dataset are not constrained to that dataset but are applicable to other datasets with similar distributions.
The importance of dataset distributional richness also plays a significant role in balancing accuracy and privacy. To achieve this, we rely on datasets that span a wide feature space, ensuring that the RFF variance captures essential data patterns. This distributional breadth is critical for extending our model’s utility across various datasets, emphasizing that conclusions are grounded in distribution characteristics rather than specific data points. As such, the robustness of our framework depends more on the diversity within a dataset’s distribution than on any individual samples, which enhances the model’s applicability to new datasets in different settings.
However, certain limitations arise when extending the DP-HD framework to larger datasets or real-world applications. While the RFF approach is computationally efficient, it may degrade when dealing with complex or non-stationary data distributions. In practice, real-world data may not fully align with assumptions of shift-invariance or smoothness required for accurate kernel approximations [83]. Additionally, the variance chosen for RFF is critical. Although our method optimizes this variance based on the data’s second moment, real-world datasets with skewed or multimodal distributions might require further refinement to capture the nuances in data patterns effectively.
The choice of exclusive versus inclusive encoding in HD also impacts the balance between accuracy and privacy. As our results indicate, encoding choices should adapt based on the distribution within each class and the distance between class centers to maintain high accuracy and privacy. Importantly, our findings reveal that the optimal standard deviation of random basis vectors remains stable across different HD configurations, as long as the dataset distribution is adequately represented. This stability further reinforces that our framework’s reliability does not depend on specific HD model parameters or sample sizes, enhancing its robustness.
We also observe that DP-HD’s uncertainty is minimal. By selecting an appropriate range of standard deviations for random basis vectors, we maintain high accuracy across privacy levels. Furthermore, in terms of inference privacy, selectively dropping dimensions from query hypervectors effectively protects query-specific information, thus misleading potential adversaries attempting to reconstruct the original data.
Finally, we benchmarked DP-HD against state-of-the-art models for both training and inference privacy. Our results indicate that DP-HD consistently outperforms these models, offering higher levels of privacy and accuracy along with faster inference times, which is essential for real-time applications such as in-situ sensing in AM.
@@ 7 Conclusion and Future Works
This paper introduced the Differential Privacy-Hyperdimensional Computing (DP-HD) framework as an innovative solution for addressing privacy concerns in Additive Manufacturing (AM). By integrating Differential Privacy (DP) mechanisms within the Hyperdimensional Computing (HD) model, the DP-HD framework ensures the protection of sensitive data during training and inference while maintaining high model performance. The key advantage of DP-HD is its explainability, enabling the prediction of noise impact on model accuracy, which is critical for in-situ sensing applications in AM. Our experiments demonstrated that DP-HD preserves privacy without significantly compromising accuracy, showcasing its robustness and suitability for real-time monitoring in AM. The ability to balance privacy and performance through explainable insights sets DP-HD apart, making it a valuable tool for enhancing manufacturing precision while safeguarding data privacy. While the DP-HD framework demonstrates significant promise, certain aspects warrant further exploration to enhance its applicability. Currently, finding the optimal encoding strategy relies on selecting training and validation subsets. Future work focuses on developing mathematical methods to automatically determine the best Standard Deviation (SD) for generating random basis vectors based on data distribution. Additionally, due to DP-HD’s success in privacy-preserving techniques, the framework can be integrated within a federated learning structure to leverage distributed data processing while maintaining robust privacy measures. Future work can explore applying the DP-HD framework in other sensitive domains, such as healthcare, financial services, and government data management, where balancing privacy and accuracy is essential for secure and effective machine learning applications.
@@ Acknowledgment
This work was supported by the National Science Foundation [grant numbers 2127780, 2312517]; the Semiconductor Research Corporation (SRC); the Office of Naval Research [grant numbers N00014-21-1-2225, N00014-22-1-2067]; the Air Force Office of Scientific Research [grant number FA9550-22-1-0253]; UConn Startup Funding, and generous gifts from Xilinx and Cisco. The authors gratefully acknowledge the valuable contributions of the National Institute of Standards and Technology (NIST), particularly Dr. Brandon Lane, who provided data for this research.
Supplementary Materials
The code used in this study is openly available for reproducibility and can be accessed at the following link: https://github.com/FardinJalilPiran/Differential-Privacy-Hyperdimensional-Computing.git.